OpenStudy (anonymous):

Need help with an equation. Hoping for someone to explain to me how to do it.

OpenStudy (anonymous):

\[\sqrt{X^2-X-5}=5\]

OpenStudy (anonymous):

Okay, so you want us to solve it.

OpenStudy (anonymous):

Square each side.

OpenStudy (anonymous):

so you get x^2-x-5=25

OpenStudy (anonymous):

Then you subtract five from both sides.

OpenStudy (anonymous):

x^2-x=20.

OpenStudy (anonymous):

That's it I guess.

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

Plug in, 5^2-5=25-5=20

OpenStudy (anonymous):

Now shower me with medals or penguin friends.

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

Thank you BUT!

OpenStudy (anonymous):

What.

OpenStudy (anonymous):

What else do you need.

OpenStudy (anonymous):

I have to pick from A:-5 B:-6 C: 6 D:-5,6

OpenStudy (anonymous):

Solve for X

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

x^2-x=20

OpenStudy (anonymous):

I would go with -5

OpenStudy (anonymous):

And 6

OpenStudy (anonymous):

So D.

OpenStudy (anonymous):

okay thanks